OpenStudy (anonymous):

Show that no point on the graph of X^2-3XY+Y^2=1 has a horizontal tangent line.

OpenStudy (anonymous):

Are you still online? I just got an answer to your question and want to know you're there before I spend several minutes typing it out. :P

OpenStudy (anonymous):

I would greatly appreciate it! =D

OpenStudy (anonymous):

Use implicit differentiation (or chain rule)\[2x-3(y+x(dy/dx))+2y(dy/dx)=0\]Expand brackets\[2x-3y-3x(dy/dx)+2y(dy/dx)=0\]Factorise and rearrange\[(2y-3x)(dy/dx)=3y-2x\]And finally\[dy/dx=(3y-2x)/(2y-3x)\]We have horizontal tangents when the derivate is equal to zero, so\[(3y-2x)/(2y-3x)=0\]\[3y-2x=0\]As the bottom part doesn't effect whether it's zero or not (continuing on next post)

OpenStudy (anonymous):

Therefore\[3y=2x\]\[y=2x/3\]Substitute this back into the original equation to get\[x^2-3x(2x/3)+(2x/3)^2=1\]\[x^2-2x^2+(4x^2)/9=1\]\[(-5/9)x^2=1\]\[x^2=-9/5\]\[x=\pm \sqrt{-9/5}\]As you may no, there are no real solutions to the square root of a negative value. Therefore there are no real x values for the derivative to be zero. Hence there are no horizontal tangents.

OpenStudy (anonymous):

know :P

OpenStudy (anonymous):

thanks so much, I was so lost lol